# Frostman random variables, entropy inequalities, and applications

### v4

@@ 4.2 Diagonal polynomials of higher degrees with independent random variables
This quantity controls both the scale and inverse scale of $\rho$.
@@
$(\Delta_{z}\rho)(x)=\rho(x+z)-\rho(z)=\sum\limits_{j=0}^{d-1}\left(\sum\limits_{k=j+1}^{d}{k\choose j}\alpha_{k}z^{k-j}\right)x^{j},$
which is a polynomial of degree d1d-1. Also assume that
@@
(i) The set ${\mathbb{R}}\setminus\mathcal{R}(\rho;\eta)$ is a union of at most $d$ open intervals. Each has radius at most $2\eta$.
@@
Let $\rho$ be a real polynomial of degree $d\geq 2$. The its derivative satisfies
$\log\mathfrak{c}(\rho^{\prime})=\log\mathfrak{c}(\rho)+O_{d}(1)$
@@
$\log\mathfrak{c}(\Delta_{z}\rho)=\log\mathfrak{c}(\rho)+O\big(\big|\log|z|\big|\big)+O_{d}(1).$
We will apply the following corollary, which is a direct consequence of Lemmas 4.4 and 4.5.
Corollary 4.7.
Let $\rho$ be a real polynomial of degree $d\geq 2$.
(i) The set ${\mathbb{R}}\setminus\mathcal{R}(\rho^{\prime};\eta)$ is a union of at most d1d-1 open intervals. Each has length at most $2\eta$. And $\rho$ is monotonic on each of the intervals.
(ii) Let $0<\eta<1$. Then for any $x\in\mathcal{R}(\rho;\eta)$,
@@
$O(\log\mathfrak{K})=O_{d,s}\big(\log_{+}C_{1}+\log_{+}C_{2}+1\big)+O_{d}\big(\log\mathfrak{c}(\rho_{1})+\log\mathfrak{c}(\rho_{2})+\log_{+}c\big).$
@@
We use induction on the degree $d$. For $d=2$, the conclusion follows from Lemma 4.3. Assume that the conclusion holds for d1d-1, with $d\geq 3$. Let $\epsilon_{1}^{\prime}=\epsilon_{1}^{\prime\prime}(d-1,s)$ be the constant obtained in the inductive hypothesis. Without loss of generality, we assume that $0<\epsilon_{1}^{\prime}<1$. For $i=1,2$, let $\eta_{i}=\min\Big\{\big(\frac{\epsilon_{1}}{128dC_{i}}\big)^{1/s_{i}},\,\frac{1}{10}\Big\}$, which satisfies $dC_{i}\eta_{i}^{s_{i}}\leq\epsilon_{1}^{\prime}/128\leq 1/4$ and $\log(1/\eta_{i})=O_{s}(\log_{+}C_{i})+O_{d,s}(1)$.
Let $E$ be the event that $X\in\mathcal{R}(\rho_{1}^{\prime};\eta_{1})$, $Y\in\mathcal{R}(\rho_{2}^{\prime};\eta_{2})$. On $E$, both $X$ and $Y$ avoid the $\eta$-neighborhoods of the critical points of $\rho_{1}$ and $\rho_{2}$, ensuring that $\rho_{1}^{\prime}(X)$ and $\rho_{2}^{\prime}(Y)$ are bounded away from zero. By Lemma 4.4(i), the complement ${\mathbb{R}}\setminus\mathcal{R}(\rho_{i}^{\prime};\eta_{i})$ is a union of d1d-1 intervals, each of length $2\eta_{i}$. Since $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, we have
@@
$\geq 1-(d-1)C_{1}\eta_{1}^{s_{1}}-(d-1)C_{2}\eta_{2}^{s_{2}}\geq 1-\epsilon_{1}/64\geq 1/2.$
We denote the conditional probability given $E$ by $P_{E}$. Let $(X_{0},Y_{0})$ be the random variable such that $(X_{0},Y_{0})\equiv(X,Y|E)$ (see Lemma A.2). One can check that $X_{0}$ and $Y_{0}$ are still independent. And $(X_{0},Y_{0})$ is marginally $(s_{1},O(C_{1});s_{2},O(C_{2}))$-Frostman. By Lemma 2.32(i),
@@
By (7) and Lemma 4.4, one has
$H_{n}(X+Y)\leq n+\log_{+}c+O(1)=n+O(\log\mathfrak{K}),$
$H_{n}(\rho_{1}(X)+\rho_{2}(Y))\leq n+\log\mathfrak{c}(\rho)+O_{d}(\log_{+}c+1)=n+O(\log\mathfrak{K}).$
Together with Lemma A.2, one obtains
$H_{n}(X_{0}+Y_{0})\leq(1+\epsilon_{1}^{\prime}/32)H_{n}(X+Y)\leq H_{n}(X+Y)+n\epsilon_{1}^{\prime}/32+O(\log\mathfrak{K}),$ (50)
and
$\begin{split}H_{n}(\rho_{1}(X_{0})+\rho_{2}(Y_{0}))&\leq(1+\epsilon_{1}^{\prime}/32)H_{n}(\rho_{1}(X)+\rho_{2}(Y))\\ &\leq H_{n}(\rho_{1}(X)+\rho_{2}(Y))+n\epsilon_{1}^{\prime}/32+O(\log\mathfrak{K}),\end{split}$ (51)
where we have used the estimate $P(E)^{-1}\leq(1-\epsilon_{1}^{\prime}/64)^{-1}\leq 1+\epsilon_{1}^{\prime}/32$.
Note that $range_{+}(X_{0})\subseteq\mathcal{R}(\rho_{1}^{\prime};\eta)\cap[-c,c]$ and $range_{+}(Y_{0})\subseteq\mathcal{R}(\rho_{2}^{\prime};\eta)\cap[-c,c]$. By Lemmas 2.13, and Corollary 4.7, we see that $\rho_{1}$ and $\rho_{2}$ are weakly $(\kappa,d-1)$-bi-Lipschitz on $range_{+}(X_{0})$ and $range_{+}(Y_{0})$, respectively, with $\log\kappa=O(\log\mathfrak{K})$.
Now, let $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ be two i.i.d. copies of $(X_{0},Y_{0})$. In particular, $X_{1},X_{2},Y_{1},Y_{2}$ are independent. Combining Lemma 2.29 and (49),
@@
$\qquad\leq 7H_{n}(\rho_{1}(X_{0})+\rho_{2}(Y_{0}))-3H_{n}(X_{0})-3H_{n}(Y_{0})+O(\log\mathfrak{K})$
$\qquad\leq 7H_{n}(\rho_{1}(X_{0})+\rho_{2}(Y_{0}))-6ns+O(\log\mathfrak{K}).$ (52)
@@
$H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2})|X_{1}-X_{2},Y_{1}-Y_{2})+H_{n}(X_{2}+Y_{2}|X_{1}-X_{2},Y_{1}-Y_{2})$
$\qquad\qquad\qquad\leq H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2}))+H_{n}(X_{2}+Y_{2}).$ (53)
@@
for any random variable depending on $(X_{1},X_{2},Y_{1},Y_{2})$. We will sum over all dyadic rationals $z_{1},z_{2}$ with $P_{E}\big(D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2}\big)>0$.
Given $D_{n}(X_{1}-X_{2})=z_{1}$ and $D_{n}(Y_{1}-Y_{2})=z_{2}$, it follows by Corollary 4.7 that
$\begin{split}&\log_{+}\Big(2^{n}\big|\big(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2})\big)-\big(\rho_{1}(X_{2}+z_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{2}+z_{2})-\rho_{2}(Y_{2})\big)\big|\Big)\\ &\qquad\qquad\qquad=O_{d}(\log_{+}c)+O(\log\mathfrak{c}(\rho_{1}))+O(\log\mathfrak{c}(\rho_{2}))+O_{d}(1)=O(\log\mathfrak{K})\end{split}$ (54)
@@
The above implied constants do not depend on $(z_{1},z_{2})$. We present the proof for $X_{2}$, as the one for $Y_{2}$ can be done similarly. Indeed, by Lemmas 2.30 and 2.31, the random variable $(X_{2},X_{1}-X_{2})$ is jointly $(s_{1},s_{1},O(C_{1}))$-Frostman. Therefore, for any ball $B(x,r)$,
@@
So (56) follows.
Now, for $i=1,2$, we take $\zeta_{i}=\min\Big\{\big(\frac{\epsilon_{1}^{\prime}}{400C_{i}}\big)^{1/s_{i}},\,1/16\Big\}$, which satisfies $\log(1/\zeta_{i})=O_{s}(\log_{+}C_{i})+O_{d,s}(1)=O(\log\mathfrak{K})$. If $n<\max\{\frac{1}{s_{1}}\log\frac{400C_{1}}{\epsilon_{1}^{\prime}},\,\frac{1}{s_{2}}\log\frac{400C_{2}}{\epsilon_{1}^{\prime}},\,4\}$, then either $\log C_{1}>ns_{1}+O_{d,s}(1)$ or $\log C_{2}>ns_{2}+O_{d,s}(1)$ or $n<4$. In view of $s_{1},s_{2}\gg s$, the term $O_{s}(\log_{+}C_{1})+O_{s}(\log_{+}C_{2})+O(1)$ may represent a negative number that is less than n-n, and (48) holds automatically. In the following, we assume that $n\geq\max\{\frac{1}{s_{1}}\log\frac{400C_{1}}{\epsilon_{1}^{\prime}},\,\frac{1}{s_{2}}\log\frac{400C_{2}}{\epsilon_{1}^{\prime}},\,4\}$. As a consequence, we have $\zeta_{i}\geq 2^{-ns_{i}-4}$ $(i=1,2)$.
We first consider the case that $|z_{1}|\geq\zeta_{1}$ and $|z_{2}|\geq\zeta_{2}$. Note that $|z_{1}|,|z_{2}|\leq 2c$. In view of Lemma 4.6, the coefficients can be bounded by
$\log\mathfrak{c}(\Delta_{z_{i}}\rho_{i})=\log\mathfrak{c}(\rho_{i})+O(\log_{+}c)+O(\log(1/\zeta_{i}))+O_{d}(1)=O(\log\mathfrak{K})$ (57)
@@
$\begin{split}&H_{n}(X_{2}+Y_{2}|z_{1},z_{2})+H_{n}((\Delta_{z_{1}}\rho_{1})(X_{2})+(\Delta_{z_{2}}\rho_{2})(Y_{2})|z_{1},z_{2})\geq n(2s+\epsilon_{1}^{\prime})\\ &\qquad+O_{d}(\log\mathfrak{c}(\Delta_{z}\rho_{1})+\log\mathfrak{c}(\Delta_{z}\rho_{2}))+\log_{+}c)+O_{d,s}\big(\log_{+}\widetilde{C}_{1}+\log_{+}\widetilde{C}_{2}+1\big).\end{split}$ (58)
@@
$\begin{split}&H_{n}(X_{2}+Y_{2}|z_{1},z_{2})+H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2})|z_{1},z_{2})\geq n(2s+\epsilon_{1}^{\prime})+O(\log\mathfrak{K})\\ &\quad+O_{d,s}\big(-n(s_{1}+s_{2})-\log P(D_{n}(X_{1}-X_{2})=z_{1})-\log P(D_{n}(Y_{1}-Y_{2})=z_{2})\big).\end{split}$ (59)
For the case $|z_{1}|<\zeta_{1}$ or $|z_{1}|<\zeta_{2}$, we use the trivial bound that
@@
$\geq H_{n}(X_{2}+Y_{2}|z_{1},z_{2})\geq\frac{1}{2}\big(H_{n}(X_{2}|z_{1},z_{2})+H_{n}(Y_{2}|z_{1},z_{2})+O(1)\big)$
$\geq ns+O_{d,s}\big(\log_{+}C_{1}+\log_{+}C_{2}-n(s_{1}+s_{2})$
$\qquad\qquad\qquad-\log P(D_{n}(X_{1}-X_{2})=z_{1})-\log P(D_{n}(Y_{1}-Y_{2})=z_{2})\big)+O(1),$ (60)
which reslults from conditional independence, (9) and Lemma 2.32.
We now average over all possible values of the discretized differences. Multiplying (59) and (60) by $P(D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2})$ and summing over all $(z_{1},z_{2})$, we obtain weighted sums of conditional entropies. The left-hand side gives
$\sum\limits_{z_{1},z_{2}}P(D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2})H_{n}(X_{2}+Y_{2}|z_{1},z_{2})=H_{n}(X_{2}+Y_{2}|X_{1}-X_{2},Y_{1}-Y_{2})$
@@
$-\sum\limits_{z_{1},z_{2}}P(D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2})\log P(D_{n}(X_{1}-X_{2})=z_{1})$
$\qquad=-\sum\limits_{z_{1}}P(D_{n}(X_{1}-X_{2})=z_{1})\log P(D_{n}(X_{1}-X_{2})=z_{1})=H_{n}(X_{1}-X_{2}).$

### v5

@@ 4.2 Diagonal polynomials of higher degrees with independent random variables
$(\Delta_{z}\rho)(x)=\rho(x+z)-\rho(x)=\sum\limits_{j=0}^{d-1}\left(\sum\limits_{k=j+1}^{d}{k\choose j}\alpha_{k}z^{k-j}\right)x^{j},$
which is a polynomial in $x$ of degree d1d-1. Also assume that
@@
(i) The set ${\mathbb{R}}\setminus\mathcal{R}(\rho;\eta)$ is a union of at most $d$ open intervals. Each has length at most $2\eta$.
@@
Let $\rho$ be a real polynomial of degree $d\geq 2$. Then its derivative satisfies
$\log\mathfrak{c}(\rho^{\prime})\leq\log\mathfrak{c}(\rho)+O_{d}(1).$
@@
$\log\mathfrak{c}(\Delta_{z}\rho)\leq\log\mathfrak{c}(\rho)+O_{d}\big(\big|\log|z|\big|+1\big).$
The following lemma is important for our application.
Lemma 4.7.
Let $\rho$ be a real polynomial of degree $d\geq 2$. Suppose that $0<\eta<1$ and $c>0$.
(i) The set ${\mathbb{R}}\setminus\mathcal{R}(\rho^{\prime};\eta)$ is a union of at most d1d-1 open intervals. Each has length at most $2\eta$.
(ii) The set $\mathcal{R}(\rho^{\prime};\eta)\cap[-c,c]$, if non-empty, is a union of at most $d$ closed intervals (including singleton set). And $\rho$ is monotonic on each of the intervals.
(iii) For any $x\in\mathcal{R}(\rho^{\prime};\eta)$,
@@
(v) The function $\rho$ is weakly $(\kappa,d)$-bi-Lipschitz on the region $\mathcal{R}(\rho^{\prime};\eta)\cap[-c,c]$, with
$\log\kappa\leq\log\mathfrak{c}(\rho)+O_{d}\big(\log_{+}c+\log(1/\eta)+1\big),$
if this region is non-empty.
Proof.
The statements (i)-(iv) are direct consequences of Lemmas 4.4 and 4.5. The statement (v) follows by combining (ii)-(iv) and Lemma 2.13. ∎
@@
$\log\mathfrak{K}=O_{d,s}\big(\log_{+}C_{1}+\log_{+}C_{2}+1\big)+O_{d}\big(\log\mathfrak{c}(\rho_{1})+\log\mathfrak{c}(\rho_{2})+\log_{+}c\big).$
@@
We use induction on the degree $d$. For $d=2$, the conclusion follows from Lemma 4.3. Assume that the conclusion holds for d1d-1, with $d\geq 3$. Let $\epsilon_{1}^{\prime}=\epsilon_{1}^{\prime\prime}(d-1,s)$ be the constant obtained in the inductive hypothesis. Without loss of generality, we assume that $0<\epsilon_{1}^{\prime}<1$.
To quantify the entropy gained beyond the trivial amount $2ns$, we write the entropy increment as
$\Psi:=H_{n}(X+Y)+H_{n}\big(\rho_{1}(X)+\rho_{2}(Y)\big)-2ns.$ (52)
Our goal is to show $\Psi\geq n\epsilon_{1}^{\prime\prime}+O(\log\mathfrak{K})$ for some $\epsilon_{1}^{\prime\prime}=\epsilon_{1}^{\prime\prime}(d,s)>0$.
Let $\mathfrak{r}=\mathfrak{r}(d,s)\geq 1$ be a parameter to be determined later. For $i=1,2$, let $\eta_{i}=\min\Big\{\big(\frac{\epsilon_{1}^{\prime}}{4\mathfrak{r}dC_{i}}\big)^{1/s_{i}},\,\frac{1}{10}\Big\}$, which satisfies $dC_{i}\eta_{i}^{s_{i}}\leq\epsilon_{1}^{\prime}/4\mathfrak{r}\leq 1/4$ and $\log(1/\eta_{i})=O_{s}(\log_{+}C_{i})+O_{d,s}(1)=O(\log\mathfrak{K})$. Let $E$ be the event that $X\in\mathcal{R}(\rho_{1}^{\prime};\eta_{1})$, $Y\in\mathcal{R}(\rho_{2}^{\prime};\eta_{2})$. By Lemma 4.7(i), the complement ${\mathbb{R}}\setminus\mathcal{R}(\rho_{i}^{\prime};\eta_{i})$ is a union of at most d1d-1 intervals, each of length at most $2\eta_{i}$. Since $(X,Y)$ is marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, we have
@@
$\geq 1-(d-1)C_{1}\eta_{1}^{s_{1}}-(d-1)C_{2}\eta_{2}^{s_{2}}\geq 1-\epsilon_{1}^{\prime}/2\mathfrak{r}\geq 1/2.$
We denote the conditional probability given $E$ by $P_{E}$. Let $(X_{0},Y_{0})$ be the random variable such that $(X_{0},Y_{0})\equiv(X,Y|E)$ (see Lemma A.2). Combining (52), one obtains
$H_{n}(X_{0}+Y_{0})+H_{n}(\rho_{1}(X_{0})+\rho_{2}(Y_{0}))$
$\qquad\leq\big(1+O(\epsilon_{1}^{\prime}/\mathfrak{r})\big)\big(H_{n}(X+Y)+H_{n}(\rho_{1}(X)+\rho_{2}(Y))\big)$
$\qquad\leq 2ns+O(n\epsilon_{1}^{\prime}/\mathfrak{r}+\Psi),$ (53)
where we have used the estimate $P(E)^{-1}\leq(1-\epsilon_{1}^{\prime}/2\mathfrak{r})^{-1}\leq 1+O(\epsilon_{1}^{\prime}/\mathfrak{r})$.
One can also check that $X_{0}$ and $Y_{0}$ are still independent. And $(X_{0},Y_{0})$ is marginally $(s_{1},O(C_{1});s_{2},O(C_{2}))$-Frostman. In view of Lemma 2.32(i),
@@
It follows from (10) that
$H_{n}(X_{0}+Y_{0})\geq\frac{H_{n}(X_{0})+H_{n}(Y_{0})}{2}+O(1)\geq ns+O(\log\mathfrak{K}).$ (55)
Note that $range_{+}(X_{0})\subseteq\mathcal{R}(\rho_{1}^{\prime};\eta)\cap[-c,c]$ and $range_{+}(Y_{0})\subseteq\mathcal{R}(\rho_{2}^{\prime};\eta)\cap[-c,c]$. By Lemma 4.7(v), we see that $\rho_{1}$ and $\rho_{2}$ are weakly $(\kappa,d)$-bi-Lipschitz on $range_{+}(X_{0})$ and $range_{+}(Y_{0})$, respectively, with $\log\kappa=O(\log\mathfrak{K})$. Then we also have
$H_{n}(\rho_{1}(X_{0})+\rho_{2}(Y_{0}))\geq\frac{H_{n}(\rho_{1}(X_{0}))+H_{n}(\rho_{2}(Y_{0}))}{2}+O(1)$
$\qquad=\frac{H_{n}(X_{0})+H_{n}(Y_{0})}{2}+O(\log\mathfrak{K})\geq ns+O(\log\mathfrak{K}).$ (56)
Combining (53), (55) and (56), one obtains
$H_{n}(X_{0}+Y_{0}),\,H_{n}(\rho_{1}(X_{0})+\rho_{2}(Y_{0}))\,\leq ns+O(n\epsilon_{1}^{\prime}/\mathfrak{r}+\Psi+\log\mathfrak{K}).$ (57)
Now, let $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ be two i.i.d. copies of $(X_{0},Y_{0})$. In particular, $X_{1},X_{2},Y_{1},Y_{2}$ are independent. By Lemma 2.21,
$H_{n}(X_{1}-X_{2})$ $\leq H_{n}(X_{1}+Y_{1})+H_{n}(X_{2}+Y_{1})-H_{n}(Y_{1})+O(1)$
$=2H_{n}(X_{0}+Y_{0})-H_{n}(Y_{0})+O(1).$
Similarly, one has $H_{n}(Y_{1}-Y_{2})\leq 2H_{n}(X_{0}+Y_{0})-H(X_{0})+O(1)$. Combining these two inequalities together with (54) and (57), we get
$H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})\leq 2ns+O(n\epsilon_{1}^{\prime}/\mathfrak{r}+\Psi+\log\mathfrak{K}).$ (58)
Combining Lemma 2.29, (54) and (55), one sees that
@@
$\qquad\leq 7H_{n}(\rho_{1}(X_{0})+\rho_{2}(Y_{0}))-3H_{n}(X_{0})-3H_{n}(Y_{0})+O(\log\kappa+\log d)$
$\qquad\leq ns+O(n\epsilon_{1}^{\prime}/\mathfrak{r}+\Psi+\log\mathfrak{K}).$ (59)
@@
$\begin{split}&H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2})|X_{1}-X_{2},Y_{1}-Y_{2})+H_{n}(X_{2}+Y_{2}|X_{1}-X_{2},Y_{1}-Y_{2})\\ &\,\,\leq H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2}))+H_{n}(X_{2}+Y_{2})\leq 2ns+O(n\epsilon_{1}^{\prime}/\mathfrak{r}+\Psi+\log\mathfrak{K}),\end{split}$ (60)
which follows from Lemma 2.4(iv), (57) and (4.2).
@@
for any random variable that is a function of $(X_{1},X_{2},Y_{1},Y_{2})$. We will sum over all dyadic rationals $z_{1},z_{2}$ with $P_{E}\big(D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2}\big)>0$. In particular, one has $z_{1},z_{2}\in[-2c-1,2c+1]$.
Given $D_{n}(X_{1}-X_{2})=z_{1}$ and $D_{n}(Y_{1}-Y_{2})=z_{2}$, it follows by Lemma 4.7(iv) that
$\log_{+}\Big(2^{n}\cdot\big|\big(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2})\big)-\big(\rho_{1}(X_{2}+z_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{2}+z_{2})-\rho_{2}(Y_{2})\big)\big|\Big)$
$\qquad\qquad\qquad=O_{d}(\log_{+}c)+O(\log\mathfrak{c}(\rho_{1}))+O(\log\mathfrak{c}(\rho_{2}))+O_{d}(1)=O(\log\mathfrak{K})$ (61)
@@
The above implied constants do not depend on $(z_{1},z_{2})$. We present the proof for $X_{2}$, as the one for $Y_{2}$ can be done similarly. Indeed, by Lemmas 2.30 and 2.31, the random variable $(X_{2},X_{1}-X_{2})$ is jointly $(s_{1},s_{1},O(C_{1}^{2}))$-Frostman. Therefore, for any ball $B(x,r)$,
@@
So (63) follows. We also mention that $\widetilde{C}_{i}\gg C_{i}$ $(i=1,2)$.
Now, for $i=1,2$, we take $\zeta_{i}=\min\Big\{\big(\frac{\epsilon_{1}^{\prime}}{\mathfrak{r}C_{i}}\big)^{1/s_{i}},\,1/16\Big\}$, which satisfies $\log(1/\zeta_{i})=O_{s}(\log_{+}C_{i})+O_{d,s}(1)=O(\log\mathfrak{K})$. If $n<\max\{\frac{1}{s_{1}}\log\frac{\mathfrak{r}C_{1}}{\epsilon_{1}^{\prime}},\,\frac{1}{s_{2}}\log\frac{\mathfrak{r}C_{2}}{\epsilon_{1}^{\prime}},\,4\}$, then either $\log C_{1}>ns_{1}+O_{d,s}(1)$ or $\log C_{2}>ns_{2}+O_{d,s}(1)$ or $n<4$. In view of $s_{1},s_{2}\gg s$, the term $O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2}+1)$ may represent a negative number that is less than 10n-10\,n, and (51) holds automatically. In the following, we assume that $n\geq\max\{\frac{1}{s_{1}}\log\frac{\mathfrak{r}C_{1}}{\epsilon_{1}^{\prime}},\,\frac{1}{s_{2}}\log\frac{\mathfrak{r}C_{2}}{\epsilon_{1}^{\prime}},\,4\}$. As a consequence, we have $\zeta_{i}\geq 2^{-n-4}$ $(i=1,2)$.
We first consider the case that $|z_{1}|\geq\zeta_{1}$ and $|z_{2}|\geq\zeta_{2}$. Recall that $|z_{1}|,|z_{2}|\leq 2c+1$. In view of Lemma 4.6, the coefficients can be bounded by
$\log\mathfrak{c}(\Delta_{z_{i}}\rho_{i})\leq\log\mathfrak{c}(\rho_{i})+O_{d}\big(\log_{+}c+\log(1/\zeta_{i})+1\big)=O(\log\mathfrak{K})$ (64)
@@
$\begin{split}&H_{n}(X_{2}+Y_{2}|z_{1},z_{2})+H_{n}\big((\Delta_{z_{1}}\rho_{1})(X_{2})+(\Delta_{z_{2}}\rho_{2})(Y_{2})|z_{1},z_{2}\big)\geq n(2s+\epsilon_{1}^{\prime})\\ &\qquad+O_{d,s}\big(\log_{+}\widetilde{C}_{1}+\log_{+}\widetilde{C}_{2}+1\big)+O_{d}\big(\log\mathfrak{c}(\Delta_{z}\rho_{1})+\log\mathfrak{c}(\Delta_{z}\rho_{2})+\log_{+}c\big).\end{split}$ (65)
@@
$\begin{split}&H_{n}(X_{2}+Y_{2}|z_{1},z_{2})+H_{n}(\rho_{1}(X_{1})-\rho_{1}(X_{2})+\rho_{2}(Y_{1})-\rho_{2}(Y_{2})|z_{1},z_{2})\geq n(2s+\epsilon_{1}^{\prime})\\ &\qquad+O(\log\mathfrak{K})+O_{d,s}\big(-2ns-\log P_{E}(D_{n}(X_{1}-X_{2})=z_{1})-\log P_{E}(D_{n}(Y_{1}-Y_{2})=z_{2})\big).\end{split}$ (66)
For the case $|z_{1}|<\zeta_{1}$ or $|z_{2}|<\zeta_{2}$, we use the trivial bound that
@@
$\geq H_{n}(X_{2}+Y_{2}|z_{1},z_{2})\geq\frac{H_{n}(X_{2}|z_{1},z_{2})+H_{n}(Y_{2}|z_{1},z_{2})}{2}+O(1)$
$\geq ns+O\big(\log_{+}C_{1}+\log_{+}C_{2}-2ns$
$\qquad\qquad-\log P_{E}(D_{n}(X_{1}-X_{2})=z_{1})-\log P_{E}(D_{n}(Y_{1}-Y_{2})=z_{2})\big)+O(1),$ (67)
which results from conditional independence, (10), (63) and Lemma 2.32.
We now average over all possible values of the discretized differences. Multiplying (66) and (67) by $P_{E}(D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2})$ and summing over all $(z_{1},z_{2})$, we obtain weighted sums of conditional entropies. The first term on the left-hand side gives
$\sum\limits_{z_{1},z_{2}}P_{E}(D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2})H_{n}(X_{2}+Y_{2}|z_{1},z_{2})=H_{n}(X_{2}+Y_{2}|X_{1}-X_{2},Y_{1}-Y_{2})$
@@
$-\sum\limits_{z_{1},z_{2}}P_{E}(D_{n}(X_{1}-X_{2})=z_{1},D_{n}(Y_{1}-Y_{2})=z_{2})\log P_{E}(D_{n}(X_{1}-X_{2})=z_{1})$
$\qquad=-\sum\limits_{z_{1}}P_{E}(D_{n}(X_{1}-X_{2})=z_{1})\log P_{E}(D_{n}(X_{1}-X_{2})=z_{1})=H_{n}(X_{1}-X_{2}).$